OpenStudy (raffle_snaffle):

You estimate that you will need $11620 at the end of year 7. How much do you have to set aside today if your bank pays 18.26% per year compounded quarterly?

OpenStudy (raffle_snaffle):

@zepdrix

OpenStudy (raffle_snaffle):

oh hey

OpenStudy (raffle_snaffle):

I have a solution...

OpenStudy (retireed):

FV = PV (1+IR/n)^n Is this the equation? I did it from memory, If so solve for PV

OpenStudy (raffle_snaffle):

we have ia = (1+ (r/m))^m-1

OpenStudy (raffle_snaffle):

That is the effective annual interest rate.

OpenStudy (retireed):

I have an answer, but I am not sure about it. I got $8501.62.

OpenStudy (raffle_snaffle):

r = nominal interest rate per year m = number of compounding subperiods per year r/m = interest rate per interest period mn = number of compounding subperiods in n years

OpenStudy (raffle_snaffle):

this monthly, quartly stuff is confusing me

OpenStudy (raffle_snaffle):

I having difficulty understanding r, ia, and m

OpenStudy (retireed):

Why m-1? Let me re-read the question.

OpenStudy (retireed):

Sorry os locked up

OpenStudy (raffle_snaffle):

we need to solve for p so this is what I did... ia = (1 + (0.06/30)^30-1 = 6.17% then I used P = F(P/F,6.17%,30) = 39,719

OpenStudy (raffle_snaffle):

what is nominal interest rate and effective interest rate?

OpenStudy (retireed):

Yes I don't know what ia could be.. r is the 18.26 / 4 M would be 7 * 4 28

OpenStudy (retireed):

My memory equation was wrong. i got my book out.

OpenStudy (raffle_snaffle):

i - r/m which is the interest rate per interest period i though and r = mn which is the number of compounding subperiods in n years

OpenStudy (raffle_snaffle):

suupose to be = not -

OpenStudy (raffle_snaffle):

since we have 18.26% quartly do I need to figure out the interest rate per year?

OpenStudy (retireed):

Compound Amount Factor (Single Payment) F = P (1+i)^N Future F Present P i ?? N number of periods

OpenStudy (raffle_snaffle):

1/4(12) = 3 so there are 3 months in one subperiod? 3+3+3+3 = 12 each 3 represents a quarter

OpenStudy (raffle_snaffle):

That is how I am interrupting this

OpenStudy (retireed):

bank pays 18.26% per year, compounded quarterly so each quarter they pay 18.26/4

OpenStudy (raffle_snaffle):

so r = 18.26%?

OpenStudy (retireed):

P = 11620 / (1+18.26/4)^(4*7)

OpenStudy (raffle_snaffle):

i = r/m = 4.56% so this is per subperiod? so every 3 months we have an interest of 4.56%

OpenStudy (retireed):

Yes in your equation r is 18.26%

OpenStudy (raffle_snaffle):

I just answer the question above myself.. lol

OpenStudy (retireed):

A mistake ..... P = 11620 / (1+0.182626/4)^(4*7) r must be in decimal format

OpenStudy (retireed):

Another mistake .... P = 11620 / (1+0.1826/4)^(4*7) r must be in decimal format

OpenStudy (raffle_snaffle):

yeah I just noticed that now... I was doing the calculation over again

OpenStudy (raffle_snaffle):

I got an answer of P = 8501.62

OpenStudy (retireed):

That's what i got from my program from my equation Now i get $3329.57

OpenStudy (retireed):

I must have did the equation wrong.

OpenStudy (raffle_snaffle):

the equataion is P = F(1+i)^-n

OpenStudy (raffle_snaffle):

I want to go back and do some previous problems. I really want to understand these problems. here is question 5 Stan and Star are buying a house for $239856. The bank is offering a nominal annual mortgage rate of 6% for 30 years, compounded monthly. What will their monthly loan payment be?

OpenStudy (raffle_snaffle):

P= ? F = 239856 n = 30

OpenStudy (retireed):

I was using N as 7*4. It should be just 7, I guess.

OpenStudy (retireed):

Go to the next question... No P is the cost of the house today P= 239856 n = 30 F = is a don't care I need a house now r = 0.06 A = the amount paid each month, the monthly loan payment

OpenStudy (raffle_snaffle):

let me solve this. I did this wrong because I inturrpted it wrong

OpenStudy (raffle_snaffle):

how do I solve this one? I need i...

OpenStudy (raffle_snaffle):

6% over 30 years would be 0.002 = 0.02% every year?

OpenStudy (retireed):

After working this problem I think the last problem N is equal to 7*4 = 28 I have bought and paid for a house. I know my answer for this one is correct. I am going back to work on the first problem. To prove that N is 28, maybe.

OpenStudy (retireed):

No, but I wish 6% per year 0.06/12 per month

OpenStudy (raffle_snaffle):

Oh i see 6% per year why not 6%/12 = 0.50 which is 50% why is it 0.06/12 = 0.005 how do you know if r = to a percentage or decimal? I am confused.

OpenStudy (retireed):

I these equations always use decimals for interest.

OpenStudy (raffle_snaffle):

so it's 0.005 = 0.05% per year over 30 years? Am I interrupting that correctly?

OpenStudy (raffle_snaffle):

so when dealing with these problems I need n because n is in therms of years. so r interest rate that is not compounded and i is interest rate that is compounded?

OpenStudy (retireed):

The answer to the first question is $3329.57 use N = 28 in your calculation Here is the excel proof amount compounded per quarter for 7 years 11620 0.1826/4 0.04565 0 $3,329.57 $151.99 1 $3,481.56 $158.93 2 $3,640.50 $166.19 3 $3,806.69 $173.78 4 $3,980.46 $181.71 5 $4,162.17 $190.00 6 $4,352.17 $198.68 7 $4,550.85 $207.75 8 $4,758.60 $217.23 9 $4,975.83 $227.15 10 $5,202.97 $237.52 11 $5,440.49 $248.36 12 $5,688.85 $259.70 13 $5,948.54 $271.55 14 $6,220.09 $283.95 15 $6,504.04 $296.91 16 $6,800.95 $310.46 17 $7,111.41 $324.64 18 $7,436.05 $339.46 19 $7,775.51 $354.95 20 $8,130.46 $371.16 21 $8,501.61 $388.10 22 $8,889.71 $405.82 23 $9,295.53 $424.34 24 $9,719.87 $443.71 25 $10,163.58 $463.97 26 $10,627.55 $485.15 27 $11,112.69 $507.29 28 $11,619.99

OpenStudy (raffle_snaffle):

7*4 = 28 why is it 28?

OpenStudy (retireed):

Sorry I was off on a tangent.. so when dealing with these problems I need n because n is in terms of years. NO N is in terms of periods if compounded monthly a period is a MONTH or 12 for a year if compounded quarterly a period is 3 months or 4 per year so r interest rate that is not compounded and i is interest rate that is compounded? r is the annual rate i is the rate for the period Digest this for a while. You were even confusing me for a second earlier.

OpenStudy (retireed):

7*4 = 28 why is it 28? 7*4 = 28 why is it 28? 7 years and 4 times a year 7*4 = 28 The first problem was compounded quarterly for 7 years.... You estimate that you will need $11620 at the end of year 7. How much do you have to set aside today if your bank pays 18.26% per year compounded quarterly?

OpenStudy (raffle_snaffle):

okay so a period in this problem is 3 months so there are 4 periods and we have 7 years so 7*4 = 28. Thank you for explaining to me what a "period" can be in terms of. That make a lot of confusion go away.

OpenStudy (retireed):

Great now in the current problem.... 30 years monthly payments How many periods N ?

OpenStudy (raffle_snaffle):

360 because every period is 1 month so we have 30*12 = 360 months

OpenStudy (retireed):

Yes!

OpenStudy (raffle_snaffle):

I am probably going to write your last statement down in my textbook somewhere so I can reference if I get confused. That helped out a lot though.

OpenStudy (raffle_snaffle):

Stan and Star are buying a house for $239856. The bank is offering a nominal annual mortgage rate of 6% for 30 years, compounded monthly. What will their monthly loan payment be? I am going to try go to figure this out.

OpenStudy (retireed):

I did it on the calculator, but I trying to find the right equation.

OpenStudy (raffle_snaffle):

SO in this case the r = 6% because it's an annual interest rate?

OpenStudy (retireed):

that is r 6% use i as 0.06 / 12 = 0.005

OpenStudy (raffle_snaffle):

kk that makes a lot of sense

OpenStudy (raffle_snaffle):

I am working on the problem

OpenStudy (raffle_snaffle):

A = 1438.05

OpenStudy (retireed):

YES! I got that answer..... Use a Series Present-Worth Factor (uniform Series) A = [ i (1+i)^N ] / [ ((1+i)^N) - 1] the - 1 is NOT part of the exponent

OpenStudy (raffle_snaffle):

yeah I used A = P[i(1+i)^n/(1+i)^n-1]

OpenStudy (retireed):

And it makes logical sense. When I first did it I used N = 30 and got a very large monthly payment.

OpenStudy (raffle_snaffle):

yeah if you use 30 you get A = 8629.75

OpenStudy (retireed):

That would be most folks, entire both paychecks for a couple.

OpenStudy (raffle_snaffle):

Cushing, Inc. invested $226821 in an account paying 7% per year, compounded monthly. How much will Cushing have in its account in 9 years?

OpenStudy (raffle_snaffle):

F = 425102.69

OpenStudy (raffle_snaffle):

@retirEEd

OpenStudy (raffle_snaffle):

For retirement planning, you decide to deposit $724 at the end of every quarter and increase your deposit by $38 each quarter . How much will you have at the end of 25 years if the bank pays a nominal annual rate of 2% compounded quarterly?

OpenStudy (retireed):

F = 425102.69 yes

OpenStudy (raffle_snaffle):

^^^working on this one now. so i've concluded that n = 100 quarters because we have 1 perirod =

OpenStudy (raffle_snaffle):

For retirement planning, you decide to deposit $724 at the end of every quarter and increase your deposit by $38 each quarter . How much will you have at the end of 25 years if the bank pays a nominal annual rate of 2% compounded quarterly?

OpenStudy (raffle_snaffle):

here I got n = 25yr*4 = 100 quarters. A = 725 and G = 38 because it's increases every quarter. I was thinking this is an arithmetic gradient series, but I don't have an euqtion for finding F knowing G, F/G.

OpenStudy (retireed):

What is the starting point? What is the ending point? There might be 99 periods since the first deposit is at the END of the first quarter.

OpenStudy (raffle_snaffle):

hold on i think i figured it out

OpenStudy (retireed):

Say you start working on Jan 1, 2000. You make your first payment on March 31, 2000 You retire 25 years later on Dec 31, 2025 and you make your last payment. I get 3 payment in the first year and 4 * 24 until Dec 31, 2025 N = 99

OpenStudy (retireed):

I don't know how to do these gradient ones on my calculator so I have use my book's equation.

OpenStudy (raffle_snaffle):

you can't do these on the calculator. you have to do them by hand. I almost got the answer.

OpenStudy (raffle_snaffle):

I got F = 316,573.70

OpenStudy (retireed):

Still working ....

OpenStudy (retireed):

I didn't get that. My number is smaller. What i did you use?

OpenStudy (raffle_snaffle):

I used P=A(P/A,i,n) + G(+/G,i,n) got P then solved for F F = P(1+I)^n

OpenStudy (raffle_snaffle):

G(P/G,i,n)*

OpenStudy (retireed):

I see one of my mistakes hold on ...

OpenStudy (retireed):

Okay I got $316,764.99 which is pretty close. There might be some round off errors in excel

OpenStudy (raffle_snaffle):

or excel is more accurate

OpenStudy (raffle_snaffle):

Well I am out of here. I have class in an hour and I need to eat something. Thanks for your help today! take care.

OpenStudy (retireed):

I always forget to add the gradient in this case $38 to the previous value plus the gradient. I was just using 725 + 38 for the entire 25 years past the first quarter.

OpenStudy (retireed):

Bye

OpenStudy (raffle_snaffle):

Oh I see

OpenStudy (raffle_snaffle):

later